OpenStudy (help):

REALLY NEED HELP PLEASE!!!! PROBLEM: If cos(t) = -3/5 where pi

OpenStudy (anonymous):

based on given info, we'll get sin t= -4/5, tan t= 4/3. cos (2t) = 2 cos^2 (t) -1 sin(2t)= 2 sin t cos t just you half angle formulae to solve cos (t/2) and sin (t/2)

OpenStudy (help):

Nope, wrong. Thanks anyways